Menlo Company distributes a single product. The companys sales and expenses for last month follow Per Unit S 20 14 Total Sales Variable expenses $318,000 222,600 Contribution margin Fixed expenses 95,400 76,800 Net operating income $ 18,600 Required: 1. What is the montnly break-even point in unit sales and in dollar sales? Break-even point in unit sales Break-even point in sales dollars units 2. Without resorting to computations, what is the total contribution margin at the break-even point? 3-a. How many units would have to be sold each month to eam a target profit of $26,400? Use the formula method 3-b. Verify your answer by preparing a contribution format income statement at the target sales level. Menlo Company Contribution Income Statement Total Per unit 4. Refer to the original data. Compute the companys margin of safety in both dollar and percentage terms. Round your percentage answer to 2 decimal places i.e.1234 should be entered as 12.34). Margin of safety 5. What is the companys CM ratio? If monthly sales increase by $70,000 and there is no change in foxed expenses, by how much would you expect monthly net operating income to increase? CM ratio Net operating income increases by
Answer #1
As per HOMEWORKLIB RULES when ther are more than 4 parts to the question then we
have to answer first 4 parts
1)Calculation of break even point in unit sales and dollar sales
Break even point= Fixed cost/ contribution per unit
Break even point= 76800/6= 12800 units
Break even point in unit sales is 12800 units
Break even point in dollar sales= unit* selling price= 12800*20
Break even point in dollar sales= $256000
2)At break even point, there is no profit no loss and the contribution margin is equal to
fixed cost.
As fixed cost is $76800 so contribution margin will also be $76800
So contribution margin is $76800
3-a) Calculation of number of units:
Number of units= (Fixed cost+ profit)/ contribution per unit
                                            =(76800+26400)/6=17200 units
Number of units= 17200 units
3-b) Contribution Format Income Statement
Particulars Total Per unit
Sales(17200*20) 3,44,000 20
Variable cost(17200*14) 2,40,800 14
Contribution margin(Sales - Variable Cost) 1,03,200 6
Fixed expenses 76,800
Net operating income 26,400
4)Calculation of margin of safety in dollar:
Margin of safety in dollars= Total sales- break even point in dollars
                                                     = 318000-256000= $62000
Margin of safety in dollars= $62000
Margin of safety in percentage= 62000/318000*100= 19.50%
Margin of safety in percentage= 19.50